# Fractional Brownian motion with mean-density interaction

### v1

@@ Abstract
Fractional Brownian motion is a Gaussian stochastic process with long-range correlations in time; it has been shown to be a useful model of anomalous diffusion. Here, we investigate the effects of mutual interactions in an ensemble of particles undergoing fractional Brownian motion. Specifically, we introduce a mean-density interaction in which each particle in the ensemble is coupled to the gradient of the total, time-integrated density produced by the entire ensemble. We report the results of extensive computer simulations for the mean-square displacements and the probability densities of particles undergoing one-dimensional fractional Brownian motion with such a mean-density interaction. We find two qualitatively different regimes, depending on the anomalous diffusion exponent $\alpha$ characterizing the fractional Gaussian noise. The motion is governed by the interactions for $\alpha<4/3$ whereas it is dominated by the fractional Gaussian noise for $\alpha>4/3$. We develop a scaling theory explaining our findings. We also discuss generalizations to higher space dimensions and nonlinear interactions as well as applications to the growth of strongly stochastic axons (e.g., serotonergic fibers) in vertebrate brains.
@@ I Introduction
For example, anomalous diffusion can be caused by slowly decaying long-time correlations between the increments (steps) of the stochastic process. The paradigmatic mathematical model for this situation is fractional Brownian motion (FBM) [13, 14], a non-Markovian self-similar Gaussian stochastic process with stationary increments. Positive, persistent correlations between the increments lead to superdiffusion ($21<\alpha<$ 2), whereas negative, anti-persistent correlations produce subdiffusion ($10<\alpha<$ 1). For $\alpha=1$, FBM is identical to normal Brownian motion with uncorrelated increments. FBM has been successfully applied to model the dynamics in a wide variety of systems including diffusion inside biological cells [15, 16, 17, 18, 19, 20], the dynamics of polymers [21, 22], electronic network traffic [23], as well as fluctuations of financial markets [24, 25].
@@
Recently, reflected FBM [26, 27, 28] was employed to explain the inhomogeneous spatial distribution of serotonergic fibers (axons) in vertebrate brains [29, 30, 31]. To this end, the set of serotonergic fibers is modeled as an ensemble of FBM trajectories that propagate inside the brain, starting from the cell bodies in the brainstem. So far, different fibers have been treated as independent in this approach. However, experimental evidence in mouse models suggests that the growth of serotonergic fibers is sensitive to the extracellular levels of serotonin (released by the fibers themselves) [32, 33, 34], and that active self-repulsion (as opposed to physical volume exclusion) contributes to the distribution of serotonergic fibers in the brain [35]. Specifically, a lack of serotonin synthesis in the developing brain increases serotonergic fiber densities in some forebrain regions [33], and a pharmacologically-induced increase in brain serotonin levels (using fluoxetine, a widely prescribed antidepressant) results in a decrease in the serotonergic fiber densities in some of the same regions [36, 37]. These findings suggest that growing fibers may be sensitive to the local coarse-grained density of the entire fiber ensemble and be repulsed from regions where this density is high.
@@
In this paper, we therefore introduce an interaction that models this idea by coupling each of the particles in a large ensemble of particles to the gradient of the total, time-integrated density of an entire ensemble. We then investigate, by means of large-scale computer simulations, the behavior of FBM under the influence of this “ mean-density” interaction. We find two qualitatively different regimes. If the anomalous diffusion exponent $\alpha$ of the underlying FBM is below 4/3, the motion is governed by the interactions whereas it is dominated by the fractional Gaussian noise for $\alpha>4/3$. To explain this interesting threshold behavior, we develop a one-parameter scaling theory.
@@
Our paper is organized as follows. We define FBM, introduce the mean-density interaction, and discuss the details of our numerical approach in Sec. II. In Sec. III, we present the simulations results for the mean-square displacement. The scaling theory is developed in Sec. IV. In Sec. V, we present simulation results for the instantaneous and integrated probability densities and compare them to the scaling theory predictions. We also consider generalizations to higher space dimensions and nonlinear interactions in Sec. VI, and we conclude in Sec. VII.
@@ II.1 Fractional Brownian motion
FBM can be defined as a continuous-time centered Gaussian stochastic process for the position $X$ of a particle that is located at the origin at time $t=0$. The covariance function of the position at later times $s$ and $t$ is given by
@@
with $\alpha$ in the range $20<\alpha<$ 2. Setting $s=t$, this relation simplifies to $\langle X^{2}\rangle=2Kt^{\alpha}$ showing that FBM leads to anomalous diffusion with anomalous diffusion exponent $\alpha$. The probability density function of the position variable takes the Gaussian form
@@
For computer simulations, it is convenient to work with a discrete-time version of FBM [38]. We discretize time by defining $x_{n}=X(t_{n})$ with $t_{n}=\epsilon n$ where $\epsilon$ is the time step and $n$ is an integer. The time evolution of the position $x_{n}$ takes the form of a random walk with identically Gaussian distributed but long-time correlated steps, governed by the recursion relation
@@
To achieve the continuum limit, the standard deviation $\sigma$ of an individual step must be small compared to the considered distances. Equivalently, the time step $\epsilon$ must be small compared to the total time. $t$. The continuum limit can thus be reached either by taking $\epsilon$ to zero at fixed $t$ or by taking $t$ to infinity at fixed $\epsilon$. In this paper, we fix $\epsilon=1$ and consider the long-time limit $t\to\infty$.
@@ II.2 Mean-density interaction
$P_{\mathrm{tot}}(x,t_{n})=\sum_{j=1}^{N}\sum_{m=1}^{n}\delta[x-x^{(j)}_{m}]~{}.$ (5)
This is an appropriate choice for the application of the process to serotonergic fibers, as discussed in Sec. I. Because each fiber is represented by an FBM trajectory, $P_{\mathrm{tot}}(x,t_{n})$ corresponds to the (coarse-grained) density of the entire ensemble of fibers grown from time $t_{0}$ to $t_{n}$. We will return to this definition and possible alterrnatives in the concluding section.
The recursion relation for the position of particle $j$,
$x_{n+1}^{(j)}=x_{n}^{(j)}+\xi_{n}^{(j)}+f(x_{n}^{(j)},t_{n})$ (6)
@@
and the force term
$f(x_{n}^{(j)},t_{n})$ $=$ $-\frac{A}{N}\left.\frac{\partial}{\partial x}P_{\mathrm{tot}}(x,t_{n})\right|_{x=x_{n}^{(j)}}$ (8)
@@
Here, $P_{\mathrm{int}}=P_{\mathrm{tot}}/N$ is the mean integrated density of the ensemble. The factor $1/N$ in the relation between the total integrated density $P_{\mathrm{tot}}$ and the force is introduced in the spirit of mean-field theory to permit a well-defined thermodynamic limit $N\to\infty$. The parameter $A$ controls the character and strength of the interaction. For positive $A$, the particles are pushed away from regions of high density, whereas they are attracted to high-density regions for negative $A$. Note that the normalization of $P_{\mathrm{int}}$ is proportional to time
@@ II.3 Simulation details
We have performed computer simulations of discrete-time one-dimensional FBM with mean-density interaction for anomalous diffusion exponents $\alpha$ in the range between 0.7 (in the subdiffusive regime) and 1.7 (deep in the superdiffusive regime). We fix the time step at $\epsilon=1$ and set $K=1/2$, leading to a variance $\sigma^{2}=1$ of the individual steps. The particles start at the origin $x=0$ at $t=0$ and perform up to $2^{27}\approx 134$ million time steps.
The correlated Gaussian random numbers $\xi_{n}$ that represent the fractional noise for each particle are precalculated before the simulation by means of the Fourier-filtering method [39]. This technique starts from a sequence of independent Gaussian random numbers $\chi_{n}$ of zero mean and unit variance (which we generate using the Box-Muller transformation with the LFSR113 random number generator proposed by L’Ecuyer [40] as well as the 2005 version of Marsaglia’s KISS [41]). The Fourier transform $\tilde{\chi}_{\omega}$ of these numbers is converted via ${\tilde{\xi}_{\omega}}=[\tilde{C}(\omega)]^{1/2}\tilde{\chi}_{\omega}$, using the Fourier transform $\tilde{C}(\omega)$ of the covariance function (4). The inverse Fourier transformation of the ${\tilde{\xi}_{\omega}}$ yields the fractional Gaussian noise.
To implement the mean-density interaction, we consider ensembles of up to $N=128$ particles. The mean integrated density $P_{\mathrm{int}}(x,t_{n})$ is collected as a histogram with a narrow bin width $\Delta x=0.1$. To achieve a smooth mean integrated density for our moderately large ensemble sizes, we replace the $\delta$ functions in the definition (5) of $P_{\mathrm{int}}(x,t_{n})$ by Gaussians of variance 0.25. The gradient in the definition of the force (9) is computed via a simple two-point formula directly from the histogram. We have confirmed that small changes of these parameters do not lead to qualitative changes of the results  [^1] If the $\delta$ functions in the definition (5) of $P_{\mathrm{int}}(x,t_{n})$ are not smoothed, however, the histogram of the mean integrated density $P_{\mathrm{int}}(x,t_{n})$ develops an unphysical oscillatory instability..
To further reduce the statistical errors, the results are averaged over up to 1080 independent ensembles.
@@ III Results: Mean-squared displacement
A detailed analysis of the data in Fig. 1 reveals two different regimes. For $\alpha=1.7$, 1.5, and 4/3, the mean-squared displacement features power-law behavior over the entire time range. Fits with $\langle x^{2}\rangle\sim t^{\bar{\alpha}}$ where both $c$ and $\bar{\alpha}$ are fit parameters yield $\bar{\alpha}$ values very close to the corresponding FBM value $\alpha$. In fact, the data can be fitted with high quality (reduced $\chi^{2}$ values below unity) with $\bar{\alpha}$ fixed at $\bar{\alpha}=\alpha$. The solid lines in Fig. 1 for $\alpha=1.7$, 1.5, and 4/3 show these fits.
@@
Figure 2: Mean-squared displacement $\langle x^{2}\rangle$ of FBM with mean-density interaction vs. time $t$ for $\alpha=0.7$ and serval values of the interaction strength $A$. The data are averages over 60 ensembles of 128 random walkers each. The resulting statistical errors are much smaller than the symbol size. The dashed line represents a fit of the long-time behavior for $A=2/5$ with $\langle x^{2}\rangle=c\,t^{4/3}$ whereas the dash-dotted line shows the FBM relation $\langle x^{2}\rangle=\sigma^{2}t^{0.7}$.
At the earliest times, the mean-squared displacements are independent of $A$ and follow the FBM relation $\langle x^{2}\rangle=\sigma^{2}t^{0.7}$. After a crossover time $t_{x}$, the behavior of the mean-squared displacement changes to $\langle x^{2}\rangle\sim t^{4/3}$ with an $A$-dependent prefactor. $t_{x}$ increases with decreasing interaction strength $A$.
This crossover behavior can be understood as follows. At short times, the integrated density $P_{\mathrm{int}}$ is small. The interaction terms (forces) (9) therefore do not yet play a role in the recursion relation (6), and the process behaves just like (non-interacting) FBM. As $P_{\mathrm{int}}$ increases with time, the interaction terms (9) also increase. At the crossover time $t_{x}$, they become comparable to the fractional Gaussian noise. The crossover time increases with decreasing interaction strength $A$ because, for smaller $A$, a larger integrated density $P_{\mathrm{int}}$ is required for the same generalized force $f$. Beyond the crossover time, the process is interaction dominated, as discussed above.

### v2

@@ Abstract
Fractional Brownian motion is a Gaussian stochastic process with long-range correlations in time; it has been shown to be a useful model of anomalous diffusion. Here, we investigate the effects of mutual interactions in an ensemble of particles undergoing fractional Brownian motion. Specifically, we introduce a mean-density interaction in which each particle in the ensemble is coupled to the gradient of the total, time-integrated density produced by the entire ensemble. We report the results of extensive computer simulations for the mean-squared displacements and the probability densities of particles undergoing one-dimensional fractional Brownian motion with such a mean-density interaction. We find two qualitatively different regimes, depending on the anomalous diffusion exponent $\alpha$ characterizing the fractional Gaussian noise. The motion is governed by the interactions for $\alpha<4/3$ whereas it is dominated by the fractional Gaussian noise for $\alpha>4/3$. We develop a scaling theory explaining our findings. We also discuss generalizations to higher space dimensions and nonlinear interactions, the relation of our process to the “true” or myopic self-avoiding walk, as well as applications to the growth of strongly stochastic axons (e.g., serotonergic fibers) in vertebrate brains.
@@ I Introduction
For example, anomalous diffusion can be caused by slowly decaying long-range correlations between the increments (steps) of the stochastic process. The paradigmatic mathematical model for this situation is fractional Brownian motion (FBM) [13, 14], a non-Markovian self-similar Gaussian stochastic process with stationary increments. Positive, persistent correlations between the increments lead to superdiffusion ($1<\alpha<2$), whereas negative, anti-persistent correlations produce subdiffusion ($0<\alpha<1$). For $\alpha=1$, FBM is identical to normal Brownian motion with uncorrelated increments. FBM has been successfully applied to model the dynamics in a wide variety of systems including diffusion inside biological cells [15, 16, 17, 18, 19, 20], the dynamics of polymers [21, 22], electronic network traffic [23], as well as fluctuations of financial markets [24, 25].
@@
Recently, reflected FBM [26, 27, 28] was employed to explain the inhomogeneous spatial distribution of serotonergic fibers (axons) in vertebrate brains [29, 30, 31]. To this end, the set of serotonergic fibers is modeled as an ensemble of FBM trajectories that propagate inside the brain, starting from the cell bodies in the brainstem. This model successfully reproduces key features of the inhomogeneous fiber density distribution in the brain (whereas normal Brownian motion would tend to produce a homogeneous distribution). So far, different fibers have been treated as independent in this approach. However, experimental evidence in mouse models suggests that the growth of serotonergic fibers is sensitive to the extracellular levels of serotonin (released by the fibers themselves) [32, 33, 34], and that active self-repulsion (as opposed to physical volume exclusion) contributes to the distribution of serotonergic fibers in the brain [35]. Specifically, a lack of serotonin synthesis in the developing brain increases serotonergic fiber densities in some forebrain regions [33], and a pharmacologically-induced increase in brain serotonin levels (using fluoxetine, a widely prescribed antidepressant) results in a decrease in the serotonergic fiber densities in some of the same regions [36, 37]. These findings suggest that growing fibers may be sensitive to the local coarse-grained density of the entire fiber ensemble and be repulsed from regions where this density is high.
@@
In this paper, we therefore introduce a stochastic process that models this idea. It consists of FBM with a “mean-density” interaction, i.e., an interaction that couples each of the particles in a large ensemble of particles to the gradient of the total, time-integrated density of an entire ensemble. We then investigate, by means of large-scale computer simulations, the behavior of this stochastic process in one dimension. We find two qualitatively different regimes. If the anomalous diffusion exponent $\alpha$ of the underlying FBM is below 4/3, the motion is governed by the interactions whereas it is dominated by the fractional Gaussian noise for $\alpha>4/3$. To explain this interesting threshold behavior, we develop a one-parameter scaling theory.
Interestingly, our process can be understood as a generalization of the “true” or myopic self-avoiding random walk [38] to fractional noise, and could therefore also be called “myopic self-avoiding fractional Brownian motion”
@@
Our paper is organized as follows. We define FBM, introduce the mean-density interaction, and discuss the details of our numerical approach in Sec. II. In Sec. III, we present the simulation results for the mean-squared displacement. The scaling theory is developed in Sec. IV. In Sec. V, we present simulation results for the instantaneous and integrated probability densities and compare them to the scaling theory predictions. Effects of fluctuations due to a finite particle number are addressed in Sec. VI. We also consider generalizations to higher space dimensions and nonlinear interactions in Sec. VII, and we conclude in Sec. VIII.
@@ II.1 Fractional Brownian motion
FBM can be defined as a continuous-time centered Gaussian stochastic process for the position $X$ of a particle that is located at the origin at time $t=0$. In the absence of boundaries, the covariance function of the position at later times $s$ and $t$ is given by
@@
with $\alpha$ in the range $0<\alpha<2$. The constant $K$, of physical dimension length2/timeα, is the generalized diffusion coefficient. Setting $s=t$, this relation simplifies to $\langle X^{2}\rangle=2Kt^{\alpha}$ showing that FBM leads to anomalous diffusion with anomalous diffusion exponent $\alpha$. The corresponding probability density function of the position variable takes the Gaussian form
@@
For computer simulations, it is convenient to work with a discrete-time version of FBM [39]. We discretize time by defining $x_{n}=X(t_{n})$ with $t_{n}=\epsilon n$ where $\epsilon$ is the time step and $n$ is an integer. The time evolution of the position $x_{n}$ takes the form of a random walk with identically Gaussian distributed but long-range correlated steps, governed by the recursion relation
@@
To achieve the continuum limit, the standard deviation $\sigma$ of an individual step must be small compared to the considered distances. Equivalently, the time step $\epsilon$ must be small compared to the total time $t$. The continuum limit can thus be reached either by taking $\epsilon$ to zero at fixed $t$ or by taking $t$ to infinity at fixed $\epsilon$. In this paper, we fix $\epsilon=1$ and consider the long-time limit $t\to\infty$.
@@ II.2 Mean-density interaction
$P_{\mathrm{tot}}(x,t_{n})=\sum_{j=1}^{N}\sum_{m=1}^{n}\delta[x-x^{(j)}_{m}]~,$ (5)
where $x^{(j)}_{m}$ is the position of particle $j$ at time step $m$, and $\delta(x)$ denotes the Dirac delta function.
This is an appropriate choice, for example, for the application of the process to serotonergic fibers, as discussed in Sec. I. In this application, each growing fiber is represented by an FBM trajectory. Fibers release serotonin along their entire length at a roughly constant rate (per time and per length of fiber). Serotonin diffuses away and decays slowly, leading to a quasi-steady state in which the local serotonin density is approximately proportional to the density of fiber segments $P_{\mathrm{tot}}(x,t_{n})$ in a given region. The growing fibers sense this density and are repulsed from high-density regions. Analogous arguments would apply to other applications in which the trajectories represent growing physical objects that persist in time.
We note that one can imagine other applications in which the influence of the density at earlier times decays with time, for example if the interaction is mediated by a chemical with a finite life time that is released by the moving particle (rather than the entire trajectory). This situation can be modeled by introducing a memory kernel in eq. (5). We will briefly consider this case in Appendix A.
In the presence of a density dependent force, the recursion relation for the position of particle $j$,
$x_{n+1}^{(j)}=x_{n}^{(j)}+\xi_{n}^{(j)}+f[x_{n}^{(j)},t_{n}]$ (6)
@@
(where $\delta_{ij}$is the Kronecker delta) and the force term
$f[x_{n}^{(j)},t_{n}]$ $=$ $-\frac{A}{N}\left.\frac{\partial}{\partial x}P_{\mathrm{tot}}(x,t_{n})\right|_{x=x_{n}^{(j)}}$ (8)
@@
[…]
@@ II.3 Simulation details
We have performed computer simulations of discrete-time one-dimensional FBM with mean-density interaction for anomalous diffusion exponents $\alpha$ in the range between 0.4 (in the subdiffusive regime) and 1.7 (deep in the superdiffusive regime). We fix the time step at $\epsilon=1$ and set $K=1/2$, leading to a variance $\sigma^{2}=1$ of the individual steps. The particles start at the origin $x=0$ at $t=0$ and perform up to $2^{27}\approx 134$ million time steps.
The correlated Gaussian random numbers $\xi_{n}$ that represent the fractional noise for each particle are precalculated before the simulation by means of the Fourier-filtering method [40]. This technique starts from a sequence of independent Gaussian random numbers $\chi_{n}$ of zero mean and unit variance (which we generate using the Box-Muller transformation with the LFSR113 random number generator proposed by L’Ecuyer [41] as well as the 2005 version of Marsaglia’s KISS [42]). The Fourier transform $\tilde{\chi}_{\omega}$ of these numbers is converted via ${\tilde{\xi}_{\omega}}=[\tilde{C}(\omega)]^{1/2}\tilde{\chi}_{\omega}$, using the Fourier transform $\tilde{C}(\omega)$ of the covariance function (4). The inverse Fourier transformation of the ${\tilde{\xi}_{\omega}}$ yields the fractional Gaussian noise  [^1] The Fourier filtering method uses a window size of $2N_{t}$ where $N_{t}$ is the number of time steps. This ensures that the covariance of the resulting increments $\xi_{n}$ is given by (4) and not distorted by the periodicity of the Fourier transformation..
To implement the mean-density interaction, we consider ensembles of up to $N=128$ particles. The mean integrated density $P_{\mathrm{int}}(x,t_{n})$ is collected as a histogram with a narrow bin width $\Delta x=0.1$. To achieve a smooth mean integrated density for our moderately large ensemble sizes, we replace the $\delta$ functions in the definition (5) of $P_{\mathrm{int}}(x,t_{n})$ by Gaussians of variance 0.25. We have confirmed that small changes of the histogram bin width and the variance of the smoothing Gaussian only lead to minuscule changes of the results  [^2] If the $\delta$ functions in the definition (5) of $P_{\mathrm{int}}(x,t_{n})$ are not smoothed, however, the histogram of the mean integrated density $P_{\mathrm{int}}(x,t_{n})$ develops an unphysical oscillatory instability.. The gradient in the definition of the force (9) is computed via a simple two-point formula directly from the histogram. To ensure the robustness of the results, we have also performed test calculations using a higher-order (4-point) gradient formula as well as calculations that do not use a histogram at all but analytically compute the derivative of the sum of the accumulated Gaussians. The latter method is restricted to times $t\lessapprox 10^{4}$ because of the high numerical effort of keeping track of all the Gaussians. All gradient algorithms give the same results (within our error bars) for the available simulation times.
To further reduce the statistical errors, the results are averaged over up to 4096 independent ensembles, depending on the parameters
@@ III Results: Mean-squared displacement
A detailed analysis of the data in Fig. 1 reveals two different regimes. For $\alpha=1.7$, 1.5, and 4/3, the mean-squared displacement features a power-law behavior over the entire time range. Fits with $\langle x^{2}\rangle=ct^{\bar{\alpha}}$ where both $c$ and $\bar{\alpha}$ are fit parameters yield $\bar{\alpha}$ values very close to the corresponding FBM value $\alpha$. In fact, the data can be fitted with high quality (reduced $\chi^{2}$ values below unity) with $\bar{\alpha}$ fixed at $\bar{\alpha}=\alpha$. The solid lines in Fig. 1 for $\alpha=1.7$, 1.5, and 4/3 show these fits.
@@
Figure 2: Mean-squared displacement $\langle x^{2}\rangle$ of FBM with mean-density interaction vs. time $t$ for $\alpha=0.7$ and several values of the interaction strength $A$. The data are averages over 60 ensembles of 128 random walkers each. The resulting statistical errors are much smaller than the symbol size. The dashed line represents a fit of the long-time behavior for $A=2/5$ with $\langle x^{2}\rangle=c\,t^{4/3}$ whereas the dash-dotted line shows the FBM relation $\langle x^{2}\rangle=\sigma^{2}t^{0.7}$. Inset: crossover time $\tau_{x}$ from FBM to interaction-dominated behavior vs. interaction strength $A$. The solid line is a power law fit, yielding an exponent of 1.05-1.05- 1.05.
At the earliest times, the mean-squared displacements are independent of $A$ and follow the FBM relation $\langle x^{2}\rangle=\sigma^{2}t^{0.7}$. After a crossover time $\tau_{x}$, the behavior of the mean-squared displacement changes to $\langle x^{2}\rangle\sim t^{4/3}$ with an $A$-dependent prefactor. $\tau_{x}$ increases with decreasing interaction strength $A$, as can be seen in the inset of Fig.  2. We have observed an analogous behavior for $\alpha=1.0$. For $\alpha=1.5$, in contrast, the data for different $A$ are essentially indistinguishable because the force terms are negligible compared to the noise.
This crossover at $\tau_{x}$ for $\alpha<4/3$ can be understood as follows. At short times, the integrated density $P_{\mathrm{int}}$ is small. The interaction terms (forces) (9) therefore do not yet play a role in the recursion relation (6), and the process behaves just like (non-interacting) FBM. As $P_{\mathrm{int}}$ increases with time, the interaction terms (9) also increase. At the crossover time $\tau_{x}$, their contribution to the displacement becomes comparable to that of the fractional Gaussian noise. The crossover time increases with decreasing interaction strength $A$ because, for smaller $A$, a larger integrated density $P_{\mathrm{int}}$ is required for the same generalized force $f$. Beyond the crossover time, the process is interaction dominated, as discussed above.
It is also interesting to visualize individual trajectories. In the noise-dominated regime $\alpha>4/3$, the trajectories closely resemble FBM trajectories because the interaction terms (9) are small compared to the fractional Gaussian noise. We therefore focus on the interaction-dominated regime, $\alpha<4/3$. Figure 3 presents trajectories for $\alpha=0.7$ with and without the mean-density interaction. For comparison, it also shows FBM trajectories with $\alpha=4/3$.
Figure 3: Plots of 16 individual trajectories, randomly chosen from simulations with $N=128$ particles. (a) FBM with mean-density interaction for $\alpha=0.7$ and $A=1/40$. (b) Non-interacting FBM, $\alpha=0.7$. (c) Non-interacting FBM, $\alpha=4/3$. (Note that FBM is self-similar; thus the appearance of the trajectories in panel (c) does not depend on the absolute scale.)
The trajectories of non-interacting FBM with $\alpha=0.7$ shown in panel (b) are highly jittery (antipersistent) because of the negative correlations of subdiffusive ($\alpha<1$) fractional Gaussian noise. This jittery motion is still visible in the presence of the mean-density interaction shown in panel (a). However, it is superposed onto a more regular motion away from the starting point (clearly visible for the trajectories furthest away from the origin (in the tails of the probability density). The comparison with panel (c) demonstrates that the trajectories of FBM with mean-density interaction do not resemble those of non-interacting FBM for $\alpha=4/3$ even though the mean-squared displacements of both processes evolve as $\langle x^{2}\rangle\sim t^{4/3}$. The $\alpha=4/3$ FBM trajectories are less jittery, even on short time scales because the fractional Gaussian noise for $\alpha>1$ is positively correlated.